OpenStudy (anonymous):

MEDALS! Erika was working on solving the exponential equation 50x = 17; however, she is not quite sure where to start. Using complete sentences, describe to Erika how to solve this equation and how solving would be different if the bases were equal.

OpenStudy (anonymous):

nErika could solve with a calculator and do trial and error until she found some power of 5 that would equal 17 (it's a little above 1.76 by this method). But being a smart person, Erika would see that the best solution would be to take logs of both sides Log 5^x = log 17....... Erika knows that log 5^x = x(log 5) x(log 5) = log 17 0.69897x = 1.2304489....divide across by 0.69897 x = 1.76055, a much more precise solution

OpenStudy (anonymous):

@Soniatonia you got that from yahoo answers...

OpenStudy (anonymous):

no I didn't, i used yahoo answers to help me solve the problem

OpenStudy (anonymous):

it's the same formula

OpenStudy (anonymous):

i have the other page open for yahoo answers and i can tell you copied and pasted it. the question was asked 1 year ago...

OpenStudy (anonymous):

Ok, if you already have the answer, why are you asking for help?

OpenStudy (anonymous):

because i want help in different words.

OpenStudy (darthvader2900):

Your welcome

OpenStudy (anonymous):

Well, i don't normally use yahoo answers for anything , but I thought that, that was what you was asking for I apologize

OpenStudy (anonymous):

it's okay. at least you tried